# Large language models can be zero-shot anomaly detectors for time series?

### v1

@@ 7 Conclusion
In this paper, we present a new challenging task to large language models. Specifically, we ask large language models to detect anomalies in time series data with no prior learning. To this end, we propose two methods Prompter, and Detector covering the prompting and forecasting paradigms, respectively. We demonstrated that LLMs can find anomalies through the forecasting paradigm (Detector method) more accurately than through prompting. We created SigLLM, a framework for converting signals into text, enabling LLMs to work with time series data. One of the major weaknesses of LLMs is their limited context window size. In SigLLM, we rely on rolling windows to chop up the time series into smaller segments. This is both inefficient and costly. Moreover, this presents a serious challenge in expanding the framework to work with multi-variate times series. With the rapid pace of advancements in LLMs, we are witnessing more models handling larger context sizes. However, they are still not at the capacity of handling time series data without segmentation. However, still with their limited capability, they are able to beat a well known transformers method and get closer to classical methods when tested against 492 signals. They fall short of the deep learning models by a factor of ×$\times 1.2$. Moreover, in anomaly detection, post-processing strategies are very critical to reveal the location of anomalies. We aim to investigate post-processing functionalities that help Prompter to filter through false alarms in future work. Similarly, a thorough exploration of error functions that bring out anomalies need to be conducted for Detector.
@@ Acknowledgments
The preparation of these instructions and the and Bib files that implement them was supported by Schlumberger Palo Alto Research, AT&T Bell Laboratories, and Morgan Kaufmann Publishers. Preparation of the Microsoft Word file was supported by IJCAI. An early version of this document was created by Shirley Jowell and Peter F. Patel-Schneider. It was subsequently modified by Jennifer Ballentine, Thomas Dean, Bernhard Nebel, Daniel Pagenstecher, Kurt Steinkraus, Toby Walsh, Carles Sierra, Marc Pujol-Gonzalez, Francisco Cruz-Mencia and Edith Elkind.
Appendix A Prompter
1import numpy as np
2
3def detect_anomalies(data, window_size, threshold):
4 moving_avg = np.convolve(data, np.ones(window_size)/window_size, mode=’valid’)
5 residual = data[window_size-1:] - moving_avg
6
7 std_dev = np.std(residual)
8 anomaly_indices = np.where(np.abs(residual) > threshold * std_dev)[0] + window_size - 1
9
10 return anomaly_indices
11
12# Example usage
13data = [...]
14window_size = 10
15threshold = 3
16anomaly_indices = detect_anomalies(data, window_size, threshold)
[…]
21
22
23
24
[…]
26
[…]
28
[…]
30
[…]
37
[…]
44
[…]
51
[…]

### v3

@@ 7 Conclusion
In this paper, we present large language models with a new and challenging task: detecting anomalies in time series data with no prior learning. To this end, we propose two methods, Prompter and Detector, covering the prompting and forecasting paradigms respectively. We demonstrate that LLMs can find anomalies through the forecasting paradigm (Detector method) more accurately than they can through the Prompter method. We present SigLLM, a framework for converting signals into text, enabling LLMs to work with time series data. A major weakness of LLMs is their limited context window size. In SigLLM, we rely on rolling windows to chop the time series up into smaller segments. This is both inefficient and costly. Moreover, it seriously challenges the possibility of expanding the framework to work with multivariate time series. As LLMs rapidly advance, more models can handle larger context sizes; however, they do not yet have the capacity to handle time series data without segmentation. Even with their limited capability, LLMs are able to beat a well-known transformer method, and to approach the performance of classical methods when tested against 492 signals. They fall short of deep learning models by a factor of ×$\times 1.2$. In anomaly detection, post-processing strategies are critical for revealing the locations of anomalies. In future work, we plan to investigate post-processing functionalities that can help Prompter filter false alarms. Similarly, we plan to conduct a thorough exploration of error functions that bring out anomalies for Detector.
@@ Acknowledgments
Our research is supported by SES S.A., Iberdrola and ScottishPower Renewables, and Hyundai Motor Company.
Change Log
Updated the acknowledgment section. Removed the previous version which had acknowledgements to the folks who had helped create the IJCAI template.